# Multi-Relational Algebra and Its Applications to Data Insights

### v3

@@ 2.1 Data Model
Slice representation of a multi-relation. In relational algebra, the unit of information processing is a tuple. In multi-relational algebra, the information unit becomes a slice. That is, operators in the multi-relational algebra are about processing slices encoded in a multi-relation. We start with the definition of slice.
@@
Definition 4 (Representing a block in a multi-relation).
@@
For (1), if there are multiple relations satisfy that condition, there are multiple ways to handle tie (for example, we can break the tie arbitrarily, or simply report an error). We leave details of handling ties later.
@@ 2.2.1 MultiSelect
Definition 7 (MultiSelect).
Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M_{1},\dots,M_{k}$ be $k$ slice transformations, where $M_{i}$ has feature set $\operatorname{{\mathcal{}F}}_{i}$, and results_predicates be predicates for filtering results. A MultiSelect is defined as
@@
slice_transformations=[$M_{1}$, ..., $M_{k}$],
results_predicates,
and is processed as follows: For each $i\in[n],j\in[k]$,
@@
    For each $\Gamma_{i}$ and $\operatorname{{\mathcal{}F}}_{j}$ (of model $M_{j}$) we compute the slice representation $\operatorname{{\mathcal{}S}}^{i}_{j}$ of $\Psi$ w.r.t. $(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
@@
    We apply $M_{j}$ to all slices in $\operatorname{{\mathcal{}S}}^{i}_{j}$, and filter the output slices using results_predicates.
@@
Let $\Psi$ be a multi-relation as specified in Figure 2, LocalTimeSeriesOutlierDetect be 1-slice transformation which detects timeseries outlier with feature set [Date, SumRevenue] (which computes a p_val for every date in a date range), results_predicates be ‘p_val < .001’, and region spaces be [Device] and [Browser]. The following MultiSelect performs outlier detection for [Date, SumRevenue] timeseries slices for each specified region space:
@@
metric_attr=SumRevenue
predicates=[p_val < .001,]
@@
  • for region space [Device], we match the relation with schema [Device, Date, SumRevenue] in $\Psi$, then we partition by Device, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
  • for region space [Browser], we match the relation with schema [Browser, Date, SumRevenue] in $\Psi$, then we partition by Browser, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and finally filter the output slice using ‘p_val < .001‘.
MultiSelect generalizes GROUP-BY GROUPING SETS and WINDOW functions. If we recognize a usual relation as a multi-relation with a single schema, then we observe that MultiSelect generalizes GROUP-BY GROUPING SETS. This is because GROUP-BY GROUPING SETS is about slice transformations that aggregate a single column. Therefore, MultiSelect also trivially generalize GROUP-BY and CUBE [GBLP96], which are special cases of GROUP-BY GROUPING SETS. MultiSelect also generalizes WINDOW functions, since we can use a slice transformation working on both the window column and metric column, and computes window-based aggregation (a more detailed explanation for this is deferred to A). Finally, we note that MultiSelect is trivially more general than all these operators since it can work on multi-relations.
@@ 2.2.2 MultiProject
The Project operator in the classic relational algebra restricts a relation to a set of attributes. In the multi-relational algebra, MultiProject restricts a multi-relation to a set of schemas. Formally, MultiProject applies the block representation procedure 4 to multiple blocks:
@@
Let $\Gamma_{1},\dots,\Gamma_{n}$ be region space schemas, and $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}$ be feature sets, then the MultiProject operator
@@
feature_sets=[$\operatorname{{\mathcal{}F}}_{1},\operatorname{{\mathcal{}F}}2,\dots,\operatorname{{\mathcal{}F}}_{m}$],
is the union of blocks for all $(\Gamma_{i},\operatorname{{\mathcal{}F}}_{j})$.
@@ 2.2.3 MultiJoin
In the classic join, we have two tuples $t,t^{\prime}$, and we have a set of columns shared by both tuples as join keys, then we merge these two tuples into one $\widetilde{t}$, subject to some conditions on the join keys. In multi-relational algebra, the situation is similar, except that we merge two slices $(r,X)$ and $(r^{\prime},X^{\prime})$ into one $(\widetilde{r},\widetilde{X})$, where $r,r^{\prime}$ act as join keys. and the merge is captured by a 2-slice transformation.
@@
slice_transformations=[$M_{1},\dots,M_{k}$],
Suppose we associate a region $r$ in $\Psi_{1}$ with another region $r^{\prime}$ in $\Psi_{2}$ that are joinable w.r.t. the specified conditions, For any 2-slice transformation $M_{i}$, it has two features sets $\operatorname{{\mathcal{}F}}_{1}$, $\operatorname{{\mathcal{}F}}_{2}$, from which we can identify a slice $S=(r,X)$ and $S^{\prime}=(r^{\prime},X^{\prime})$, and so we can compute a slice using $M(S,S^{\prime})$. We repeat this for all models, and merge all output slices into a multi-relation.
@@ 3.1 Frequent Itemset Mining
    Astute readers may argue that our approach requires encoding transactions in a very wide table, with most of the table 00 (the number of columns is the number of distinct items), and it is wasteful to do this encoding. There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section B.
@@ 5 Conclusion
Appendix A Emulating WINDOW functions
@@
For this query, we first partition the table by Department, so each partition is a table on schema [Name, Age, Salary]. Then for each partition, we sort by Age, and compute a moving average of Salary for each row. We can emulate this using MultiSelect as follows:
@@
Appendix B Encoding a sparse relation

### v4

@@ 2.1 Data Model
Slice representation of a multi-relation. In relational algebra, the unit of information processing is a tuple. In multi-relational algebra, the information unit becomes a slice. That is, operators in the multi-relational algebra are about processing slices encoded in a multi-relation. For this, a multi-relational operator will first represent a multi-relation as a set of slices. We start with the definition of slice.
@@
We now define a first slice representation method using minimal schema matching. This representation method is mostly commonly used in our current applications.
Definition 4 (Slice representation of a block via Minimal Schema Matching).
@@
A visualization of a slice representation is shown in Table 2.
Region $\operatorname{{\mathcal{}F}}$
$r_{1}$ $X_{1}$
$r_{2}$ $X_{2}$
Table 2: Slice representation of a block for a multi-relation. We have a region space $\Gamma$ and one feature set $\operatorname{{\mathcal{}F}}$. The slice representation is organized as a table where each row is a slice.
Note that for Minimal Schema Matching, it might happen that for a general multi-relation, there are multiple relations satisfy that condition. In our applications, this is usually not a concern since they mostly want to work on the results of a previous groupby grouping sets query. Note also that there are multiple ways to handle tie (for example, we can break the tie arbitrarily, or simply report an error). There are cases where we need more stringent ways than Minimal Schema Matching for slice representation, which we describe in Appendix A.1.
@@
With block representation, we now define slice representation for multiple region spaces and one feature set. In Appendix A.2, we will generalize this to multiple feature sets.
Definition 5 (Slice representation of a multi-relation with multiple region spaces and one feature set).
Let $\Gamma_{1},\dots,\Gamma_{N}$ be region spaces, and $\operatorname{{\mathcal{}F}}$ be one feature set. A slice representation of a multi-relation $\Psi$ w.r.t. the region spaces and feature sets is a table where each row is a slice. Specifically, row $i$ is of the form $(r_{i},X_{i})$, where $r_{i}$ is a region tuple of schema of some region space, and $X_{i}$ is a feature table of schema $\operatorname{{\mathcal{}F}}$
@@ 2.2.1 MultiSelect
We are ready to define MultiSelect. For the main text we will define MultiSelect for one slice transformation. Appendix B generalizes this to multiple slice transformations with different feature sets.
Definition 8 (MultiSelect with one slice transformation).
Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M$ be a slice transformation which requests feature set $\operatorname{{\mathcal{}F}}$. Finally, selection_condition is a predicate for filtering results. A MultiSelect is defined as:
@@
slice_transformations=[$M$],
selection_condition,
The evaluation of MultiSelect is as follows:
@@
    Slice representation. Given the region spaces $\Gamma_{1},\dots,\Gamma_{n}$ and feature sets $\operatorname{{\mathcal{}F}}$, we compute a slice representation $\rho$ of $\Psi$, where row $i$ of the slice representation $\rho$ is a slice $(r_{i},X_{i})$, where $X_{i}$ has schema $\operatorname{{\mathcal{}F}}$.
@@
    Slice transformation and output slice representation. For row-$i$, we evaluate each slice transformation $M$ and compute an output slice $(r^{\prime}_{i},X^{\prime}_{i})$.
  3. 3.
    Filter slice rows. The selection_condition is used to filter each output slice. (specifically, if $X_{i}^{\prime}$ is filtered as an empty relation, the slice is filtered).
  4. 4.
    Output multi-relation. We group output slices according to their schemas (the only schema non-uniformity comes from region spaces from our assumptions), and each group of slices are union into one relation. The result is a multi-relation, which serves as the output.
@@
Let $\Psi$ be a multi-relation as specified in Figure 2, LocalTimeSeriesOutlierDetect be 1-slice transformation which detects timeseries outlier with feature set [Date, SumRevenue]. Specifically, given a timeseries, it detects whether there is an outlier for a given date, and outputs a p_val for that date to indicate the strength of evidence that it is an outlier. Finally, let selection_condition be ‘p_val < .001’, and region spaces be [Device] and [Browser]. The following MultiSelect performs outlier detection for [Date, SumRevenue] timeseries slices for each specified region space:
@@
metric_attr=SumRevenue,
detection_date=’2024-09-01’,
selection_condition=[p_val < .001],
@@
  • for region space [Device], we match the relation with schema [Device, Date, SumRevenue] in $\Psi$, then we partition by Device, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and detect outliers for the date 2024-09-01. We filter the output with ‘p_val < .001‘, thus only finding devices for which the p-value on 2024-09-01 is less than .001.
  • for region space [Browser], we match the relation with schema [Browser, Date, SumRevenue] in $\Psi$, then we partition by Browser, apply LocalTimeSeriesOutlierDetect on the resulting [Date, SumRevenue], and detect outliers for the date 2024-09-01. We filter the output with ‘p_val < .001‘, thus only finding browsers for which the p-value on 2024-09-01 is less than .001.
MultiSelect generalizes GROUP-BY GROUPING SETS and WINDOW functions. If we recognize a usual relation as a multi-relation with a single schema, then we observe that MultiSelect generalizes GROUP-BY GROUPING SETS. To see this, note that each grouping set can be treated as a region space, and aggregation of several columns can be captured by a slice transformation that requests those columns as a feature set, and aggregate each column using desired aggregation functions. For example, suppose we want to do aggregations SUM(Revenue) and MAX(Clicks), then we can request a feature set [Revenue, Clicks], and the slice transformation performs sum and max aggregation on the respective column. With this, MultiSelect also generalizes CUBE [GBLP96], which are special cases of GROUP-BY GROUPING SETS. Note that in these cases, the slice transformation outputs a single tuple as output. MultiSelect also generalizes WINDOW functions, where we can use a slice transformation working on both the window column and metric column, and computes window-based aggregation, which in this case the output is a table, not a tuple. The details are deferred to Appendix C. Finally, we note that MultiSelect is trivially more general than all these operators since it can work on multi-relations.
@@ 2.2.2 MultiProject
The Project operator in the classic relational algebra restricts a relation to a set of attributes. In the multi-relational algebra, MultiProject restricts a multi-relation to a set of schemas. Formally, MultiProject applies the block representation procedure 4 to multiple blocks induced by different region spaces (similar to MultiSelect, for the main text we only consider a single feature set; and the case for multiple feature sets can be handled using similar ideas described in Appendix B)
@@
Let $\Gamma_{1},\dots,\Gamma_{n}$ be region space schemas, and $\operatorname{{\mathcal{}F}}$ be feature sets, then the MultiProject operator
@@
feature_sets=[$\operatorname{{\mathcal{}F}}$],
computes a slice representations for all blocks $\{(\Gamma_{i},\operatorname{{\mathcal{}F}})\ |\ i\in[n]\}$, and then convert the slice representation to an output multi-relation, same as described in MultiSelect (Definition 8).
@@ 2.2.3 MultiJoin
In the classic join, we have two tuples $t,t^{\prime}$, and we have a set of columns shared by both tuples as join keys, then we merge these two tuples into one $\widetilde{t}$, subject to some conditions on the join keys. In multi-relational algebra, the situation is similar, except that we merge two slices $(r,X)$ and $(r^{\prime},X^{\prime})$ into one $(\widetilde{r},\widetilde{X})$, where $r,r^{\prime}$ act as join keys. and the merge is captured by a 2-slice transformation. We formalize this below. Similar to MultiSelect and MultiProject, we only define the case for one slice transformation, which can be generalized to multiple slice transformations using ideas described in Appendix B.
@@
slice_transformations=[$M$],
where the 2-slice transformation $M$ has two features sets $\operatorname{{\mathcal{}F}}_{1}$, $\operatorname{{\mathcal{}F}}_{2}$, and $\Gamma_{1},\dots,\Gamma_{n}$ are region spaces. For evaluating this operator, we do a slice representation $\rho_{1}$ for $\{(\Gamma_{i},\operatorname{{\mathcal{}F}}_{1})\ |\ i\in[n]\}$, and another slice representation $\rho_{2}$ for $\{(\Gamma_{i},\operatorname{{\mathcal{}F}}_{2})\ |\ i\in[n]\}$. Then for a slice $\theta=(r,X)\in\rho_{1}$ and $\theta^{\prime}=(r^{\prime},X^{\prime})\in\rho_{2}$, if regions $r$ and $r^{\prime}$ are joinable w.r.t. the specified conditions, then we compute an output slice using $M(\theta,\theta^{\prime})$. The final result is an output slice representation, which we convert to a multi-relation same as in MultiSelect (Definition 8).
@@ 3.1 Frequent Itemset Mining
    Astute readers may argue that our approach requires encoding transactions in a very wide table, with most of the table 00 (the number of columns is the number of distinct items), and it is wasteful to do this encoding. There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section D.
@@ 5 Conclusion
Appendix A More on data modeling
A.1 More on slice representations
We describe two methods for more rigid slice representations. Both methods require more information from the user, passing as parameters of a multi-relational operator.
Rigid Schema Matching. In the Minimal Schema Matching method, we ask the user to pass in a list of feature sets $[\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}]$. In the Rigid Schema Matching, for each feature schema we ask the user to pass in an additional “source feature schema”, so it is like $[(\operatorname{{\mathcal{}F}}_{1},\widehat{\operatorname{{\mathcal{}F}}}_{1}),\dots,(\operatorname{{\mathcal{}F}}_{k},\widehat{\operatorname{{\mathcal{}F}}}_{k})]$. At the time of slice representation, for each region space $\Gamma_{i}$ and $(\operatorname{{\mathcal{}F}}_{j},\widehat{\operatorname{{\mathcal{}F}}}_{j})$, we will construct schema $\widehat{S}_{ij}=\Gamma_{i}\cup\widehat{\operatorname{{\mathcal{}F}}}_{j}$, and find a relation $R$ in the multi-relation with schema exactly $\widehat{S}_{ij}$. If there are multiple matching, it is an error. From $R$, then we project to $\operatorname{{\mathcal{}F}}_{j}$ for features. For readers familiar with SQL, one can think of $\widehat{F}_{j}$ as providing a FROM clause to indicate the relation we want to work on.
Minimal Containing Dimensions Schema Matching. In the Minimal Containing Dimensions Schema Matching, we ask the user to additionally pass in a list $D$ of dimension attributes (which is a subset of all attributes). Now given a region schema $\Gamma_{i}$ and feature set $\operatorname{{\mathcal{}F}}_{j}$ (let $S_{ij}=\Gamma_{i}\cup\operatorname{{\mathcal{}F}}_{j}$, the desired relation is identified as follows: We first go through all relations that contain the schema $S_{ij}$, and pick relations such that their dimension set (intersecting with $D$) is minimal. Then among all these relations, we pick the one with the minimal schema. For ties, we can either break the tie arbitrarily, or report an error.
A.2 Generalizing slice representation to multiple feature sets
Definition 17 (Slice representation of a multi-relation with multiple region spaces and multiple feature sets).
Let $\Gamma_{1},\dots,\Gamma_{N}$ be region spaces, and $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{M}$ be feature sets. A slice representation of a multi-relation $\Psi$ w.r.t. the region spaces and feature sets is a table where each row is called a slice tuple or slice row. Specifically, row $i$ is of the form $(r_{i},X^{i}_{1},X^{i}_{2},\dots,X^{i}_{M})$, where $r_{i}$ is a region tuple of schema of some region space, and $X^{i}_{j}$ is a feature table of schema $\operatorname{{\mathcal{}F}}_{j}$ ($j\in[M]$). A visualization of a slice representation is shown in Table 4.
Region $\operatorname{{\mathcal{}F}}_{1}$ $\operatorname{{\mathcal{}F}}_{2}$cdots $\operatorname{{\mathcal{}F}}_{M}$
$r_{1}$ $X^{1}_{1}$ $X^{1}_{2}$cdots $X^{1}_{M}$
$r_{2}$ $X^{2}_{1}$ $X^{2}_{2}$cdots $X^{2}_{M}$
Table 4: Slice representation of a multi-relation. We have multiple region spaces $\Gamma_{1},\dots,\Gamma_{N}$, and multiple feature sets $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{M}$. The slice representation is organized as a table of multiple slice tuples (or slice rows). Each row consists of a region tuple, followed by $M$ feature tables, one for each schema $\operatorname{{\mathcal{}F}}_{j}$, $j\in[M]$.
Appendix B Generalizing MultiSelect to multiple feature sets
B.1 Multiple tuple-output slice transformations
We now define MultiSelect with multiple tuple-output slice transformations (instead of one table-ouput slice transformation).
Definition 18 (MultiSelect with multiple tuple-output slice transformations).
Let $\Psi$ be a multi-relation, $\Gamma_{1},\dots,\Gamma_{n}$ be $n$ region space schemas, $M_{1},\dots,M_{k}$ be $k$ slice transformations, where $M_{i}$ requests feature set $\operatorname{{\mathcal{}F}}_{i}$ and is region-invariant. Further, the output feature schemas of slice transformations are pair-wise disjoint. For simplicity, we assume that on input $(r,X)$, $M_{i}$ outputs $(r,\zeta)$, where $\zeta$ is a single tuple. Finally, selection_condition is a predicate for filtering results (the specification is the same as selection condition of SELECT operator in the relational algebra; we explain this below). A MultiSelect is defined as:
MultiSelect(
$\Psi$,
region_spaces=[$\Gamma_{1}$, ..., $\Gamma_{n}$]
slice_transformations=[$M_{1}$, ..., $M_{k}$],
selection_condition,
The evaluation of MultiSelect is as follows:
  1. 1.
    Slice representation. Given the region spaces $\Gamma_{1},\dots,\Gamma_{n}$ and feature sets $\operatorname{{\mathcal{}F}}_{1},\dots,\operatorname{{\mathcal{}F}}_{k}$, we compute a slice representation $\rho$ of $\Psi$, where row $i$ of the slice representation $\rho$ is a slice tuple $(r_{i},X^{i}_{1},\dots,X^{i}_{k})$, where $X^{i}_{j}$ has schema $\operatorname{{\mathcal{}F}}_{j}$.
  2. 2.
    Slice transformation and output slice representation. For a slice row, we evaluate each slice transformation $M_{j}$ on each of the slice $(r_{i},X^{i}_{j})$, and compute output slice $(r_{i},\zeta^{i}_{j})$. The result is another slice representation, where row-$i$ is $(r_{i},\zeta^{i}_{1},\dots,\zeta^{i}_{k})$.
  3. 3.
    Filter slice rows. The selection_condition is used to filter each slice tuple of the output from the previous step. Specifically, since $r_{i}$ and $\zeta^{i}_{j}$ are all tuples, all tuples can be concatenated into one tuple, so selection_condition applies to filter each tuple. (this is why selection_condition can be specified the same way as in the relational algebra)
  4. 4.
    Output multi-relation. We group tuples according to their schemas (the only schema non-uniformity comes from region spaces from our assumptions), and each group of tuples are union into one relation. The result is a multi-relation, which serves as the output.
B.2 Multiple table-output slice transformations
[…]
In this section, we investigate the question of slice representation output where there are multiple slice transformations, and each output a feature table. In particular, we relax the requirement and allow output schemas of different slice transformations to have overlap. For simplicity, we still require that all slice transformations are region-invariant, though the argument can be carried over to allow slice transformation that are not region-invariant.
We now consider the general situation. We have an output slice row of the form $(r,Z_{1},Z_{2},\dots,Z_{k})$, where $Z_{i}$ is the output feature table of the slice transformation $M_{i}$. Further, $schema(Z_{i})\cap schema(Z_{j})$ may be non-empty. The first question is why do we want to merge these feature tables? Can’t we just have a multi-relation where we create one schema for each $(r,Z_{i})$? The first motivation is exactly as in the case join, we want to meld fragmented information together, as much as possible. Second, following the first motivation, a user may want to access two output columns, one from $Z_{i}$ and one from $Z_{j}$, and we need that there is one single relation in the multi-relation to hold both columns.
One solution to this question is to introduce a new concept, called dimension, into the theory. A dimension is an attribute which serves the role of determining whether two tables can be merged via join. Only a subset of attributes may be dimensions. Specifically, we would join two feature tables $Z_{i}$ and $Z_{j}$ if and only if they share the same dimension set. When that happens, we do a full outer-join on the dimension columns. Specifically, we say that $Z_{i}\sim Z_{j}$, if $dims(Z_{i})=dims(Z_{j})$. $\sim$ is an equivalence relation, and so we can partition $Z_{1},\dots,Z_{k}$ into equivalence classes, and within each we can join. With this, therefore we can output an output slice row $(r,Z^{\prime}_{1},\dots,Z^{\prime}_{q})$, where $q$ is the number of equivalence classes. Therefore, dimensions allow us to merge information as much as possible, as indicated by the user. To implement this, we extend the MultiSelect operator with an additional output_feature_dimensions parameter, which allows the user to configure what are the all possible output feature dimensions.
We give a concrete example of this. Suppose that we have three slice transformations. The first performs a smoothing on features [Date, Cost]. The second performs a different smoothing on features [Date, Clicks]. The third computes a CostPerClick on features [Cost, Clicks]. Suppose that the input is a multi-relation $\Psi$ with two relations, the first relation $R_{1}$ with schema [Date, Cost, Clicks], and the second $R_{2}$ with schema [Cost, Clicks]. This multi-relation may be obtained using a group-by grouping sets query, such as SELECT Date, SUM(Cost), Sum(Clicks) FROM r GROUP BY GROUPING SETS [], [Date]. Note that the region space is empty here. Therefore, a slice row looks like $(r=(),Z_{1},Z_{2},Z_{3})$ where $Z_{1}$ is a table on [Date, Cost], $Z_{2}$ is a table on [Date, Clicks], and $Z_{3}$ is a tuple (CostPerClick=XXX). A user may naturally want to join $Z_{1}$ and $Z_{2}$, since later we want to analyze Cost and Clicks together. With dimensions, the user can then configure output_feature_dimensions to be [Date]. Then we have two equivalence classes, $[Z_{1},Z_{2}]$ and $[Z_{3}]$, and we will join the tables in the first equivalence class on Date and get $Z_{1}^{\prime}$, which has schema [Date, Cost, Clicks]. The output final slice row is $(r=(),Z_{1}^{\prime},Z_{2}^{\prime}=Z_{3})$.
Appendix C Emulating WINDOW functions
@@
For this query, we first partition the table by Department, so each partition is a table on schema [Name, Age, Salary]. Then for each partition, we sort by Age, and compute a moving average of Salary for each row. We can emulate this using MultiSelect (with a single slice transformation but which outputs a table) as follows:
@@
Appendix D Encoding a sparse relation